OpenStudy (anonymous):

solve by factoring: t^3-t=0 Please explain how you did this! Thanks!

OpenStudy (anonymous):

t common factor. t(t^2-1)=0 t=0 or t^2-1=0, t^2=1, t=1. T= 0 or 1

OpenStudy (anonymous):

=t(t^2-1)=0 =t(t-1)(t+1)=0 t=0,t=1,t=-1

OpenStudy (anonymous):

*t=+-1

OpenStudy (anonymous):

so is it \[\pm1\]